Consider the following augmented matrix [12]   1 −1 2 1 3 −1 5 −2 −4 2 x 2 − 8 x + 2   . Determine the values of x for which the system has (i) no solution, (ii) exactly one solution, (iii) infinitely many solutions.


1
Expert's answer
2021-06-08T11:57:29-0400

Question

consider the following augmented matrix

"\begin{bmatrix}
 \def\arraystretch{1.5}
 \begin{array}{ ccc:}
1&-1&2&1 \\ 
3&-1&5&-2 \\ 

-4&2&x^2-8&x+2\\ 
\end{array}
\end{bmatrix}"

Determine the values of x for which the system has

(i)no solution

(ii)exactly one solution

(iii)infinitely many solutions.



solution

Here


"[A|B]=\begin{bmatrix}
 \def\arraystretch{1.5}
 \begin{array}{ ccc:}
1&-1&2&1 \\ 
3&-1&5&-2 \\ 

-4&2&x^2-8&x+2\\ 
\end{array}
\end{bmatrix}\\\&\space A=\begin{bmatrix}
 \def\arraystretch{1.5}
 \begin{array}{ ccc}
1&-1&2& \\ 
3&-1&5& \\ 

-4&2&x^2-8&\\ 
\end{array}
\end{bmatrix}"


"\begin{bmatrix}
 \def\arraystretch{1.5}
 \begin{array}{ ccc:}
1&-1&2&1 \\ 
3&-1&5&-2 \\ 

-4&2&x^2-8&x+2\\ 
\end{array}
\end{bmatrix}"


step 1:

multiply 3 to row 1 and subtract from row 2


"\begin{bmatrix}
 \def\arraystretch{1.5}
 \begin{array}{ ccc:}
1&-1&2&1 \\ 
0&2&-1&-5 \\ 

-4&2&x^2-8&x+2\\ 
\end{array}
\end{bmatrix}"


step2:

multiply 4 to row 1 and add to row 3


"\begin{bmatrix}
 \def\arraystretch{1.5}
 \begin{array}{ ccc:}
1&-1&2&1 \\ 
0&2&-1&-5 \\ 

0&-2&x^2&x+6\\ 
\end{array}
\end{bmatrix}"

step3:

row 2 add to row 3


"\begin{bmatrix}
 \def\arraystretch{1.5}
 \begin{array}{ ccc:}
1&-1&2&1 \\ 
0&2&-1&-5 \\ 

0&0&x^2-1&x+1\\ 
\end{array}
\end{bmatrix}"


now we want system has

(i)no solution

"If rank(A) < rank(A|B), then the system is no solution"


for no solution

"rank(A)=2 \implies x^2-1=0\implies x=+1 \space ,-1\\

rank(A|B)=3\implies x+1\neq0\implies x\neq-1\\"


hence both condition give value of x=+1

(ii)exactly one solution

"If rank(A) = rank(A|B) = the number of rows in x, then the system has exactly one solution. "


for exactly one solution

"rank(A) =number\space of \space rows=3\implies x^2-1\neq0\implies x\neq+1,-1\\
\\
 rank(A|B) =number\space of \space rows=3\implies x+1\neq0 \implies x\neq-1"

hence both condition give value of "x\isin R ; x\neq +1,-1"

(iii)infinitely many solutions.

If rank(A) = rank(A|B) < the number of rows in x, then the system has infinitely -many solutions.

for infinitely -many solutions.

"rank(A) = rank(A|B) < \space number \space of \space rows \space 
hence\\
rank(A) =2\implies x^2-1=0 \implies x=+1,-1\\
rank(A|B)=2\implies x+1=0 \implies x=-1"

hence both condition give value of x=-1